OpenStudy (abdela25):

explain how to determine whether two lines are parallel or perpendicular. INCLUDE ALL CASES! MEDALS REWARDED

OpenStudy (solomonzelman):

Parallel lines: have the same slope Perpendicular lines: for line with slope m, perpendicular line would have slope of -1/m

OpenStudy (anonymous):

To determine if line are parallel, perpendicular, or neither, you must compare the slopes of the lines. In the slope-intercept formula y = mx + b, the "m" is the slope. RULES: If the slopes are the SAME = parallel If the slopes are OPPOSITE RECIPROCALS = perpendicular If neither of these occur = neither parallel nor perpendicular So just compare the slopes (m) in slope-intercept form (y = mx + b)

OpenStudy (abdela25):

@mangorox @SolomonZelman thanks guys !! great help

OpenStudy (solomonzelman):

YW

OpenStudy (disco619):

To see if two lines are Perpendicular. We will see the Angle they make where they meet, it should be 90. And the product of their Slopes is -1 m1*m2 = -1 or m1 = -1/m2 Perpendicular Test:  To see if two lines are Parallel we will see their Slopes, they should be same: m1 = m2

OpenStudy (abdela25):

thanks @Disco619 added a few extra words to the other help greatly appreciated